OpenStudy (anonymous):

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies is

OpenStudy (anonymous):

total number of candies is \(6+3+2=11\) and the number of ways to pick \(3\) out of the \(11\) is \(\binom{11}{3}=\frac{11\times 10\times 9}{3\times 2}=11\times 5\times 3=165\) and that will be your denominator

OpenStudy (anonymous):

6 ways to pick one out of the 6 cherries, 1 way to pick 2 out of the 2 lemons, so your numerator will be \(6\) and your answer is therefore \[\frac{6}{165}\]

OpenStudy (anonymous):

6/165=0.0363=4%?